OpenStudy (anonymous):

Input the equation for the following line in standard form Ax + By = C. Reduce all fractional answers to lowest terms. A line that passes through the origin and is parallel to x + y = 6.

OpenStudy (aroub):

do u have any points? like (1,6)

OpenStudy (anonymous):

no :(

OpenStudy (aroub):

y=-x+0 (maybe)

OpenStudy (anonymous):

x + y = 6 write in standard forn y = -x -6 reqd equation of line is y = -x

OpenStudy (aroub):

ye bec. it passes through the origin

OpenStudy (anonymous):

y = -x + 6 is the standard form y = -x is the answer

OpenStudy (anonymous):

you are correct aroub